A canoe covers a distance of 40 miles downstream in 6 hours. Returning upstream, it takes the canoe three times as long to cover 3/5 the distance. Find the rate of the current and the rate of the canoe in still water.
Note: The answer alone is insufficient. You must not only show your work (define your variables and write two equations, solving the system) but also explain your process in order to receive full credit for you answer.

Answer by scott8148(6628)

let s=rate of stream (current) and c=rate of canoe

going downstream, the rates add ___ going upstream, the current is subtracted from the canoe
___ the calculations are in miles and hours

r = d / t

down ___ c + s = 40 / 6 = 20/3

up ___ c - s = [(3/5) * 40] / (3 * 6) = 4/3

adding the equations (eliminating s) ___ 2c = 24/3 ___ c = 4 mph

substituting (to find s) ___ (4) + s = 20/3 ___ s = 20/3 - 12/3 = 8/3 mph